Run some fidicual scans for the apparent frequency switch problem
total power, no cal in mode 4.
total power, cal, no doppler in mode 4
frequency switch, no doppler in mode 4
total power, cal, doppler in mode 4
frequency switch, doppler in mode 4.
Repeat for mode 24
Run it as fast as possible.
Do at L-band with noise source on one polarization. Running noise source on four banks.
Let's also try one with only doppler tracking.
Jason hooked it up so that odd numbers have noise source.
3 scans of 60s.
Scan 33-36: tsys_issue_mode4_nocal_nofsw_nodoppler. tint=0.01 (max for mode).
Scan 37-39: tsys_issue_mode4_cal_nofsw_nodoppler. tint=0.028 (max for mode with noise diode, no doppler). looks like bary was set in the config, but I'm slewing to 3C147 and lo1 doesn't appear to be updating, so we aren't doppler tracking. Might want to re-run.
Scan 40: tsys_issue_mode4_cal_fsw_nodoppler. tint=0.32s (max for mode with noise diode, fsw, no doppler). had keywords wrong. Trying again